A projectile Is projected vertically upwards with an initial velocity of 40m/s.two seconds later a second projectile is also projected vertically from the same point with the same initial velocity.

calculate:

a)the maximum height reached by each body

b)the time taken to reach maximum height

c)the time taken for each projectile to return to the point of projectile

d)the time taken and distance moved by each projectile when they pass each other.




1
Expert's answer
2021-07-05T17:59:15-0400

a)


"H=\frac{40^2}{2(9.8)}=82\ m"

b)


"t=\frac{40}{9.8}=4.1\ s"

c)


"T=2(4.1)=8.2\ s"

d)


"40t-0.5(9.8)t^2=40(t-2)-0.5(9.8)(t-2)^2\\t=5.1\ s"

"d_1=82+0.5(9.8)(1.1)^2=88\ m"

"d_2=2(82)-88=76\ m"